# Leveraging Synthetic Data for Question Answering with Multilingual LLMs in the Agricultural Domain

### v1

@@ Appendix A Tested Prompts
Appendix B Prompt Intuition
The following summarizes the design intuition behind the six tested prompts used for agricultural QA generation:
@@ Appendix B Prompt Intuition
  • Prompt 2: This prompt was designed to simulate a highly knowledgeable agricultural expert, embedding deep domain-specific authority into the model’s persona. The aim was to guide the model to generate precise, well-informed, and context-aware answers like domain experts.
  • Prompt 3: This prompt was designed to extract detailed, accurate, and context-sensitive responses by positioning the model as a multilingual agricultural expert. Emphasis was placed on reasoning, cultural relevance, and linguistic alignment to ensure comprehensive, well-researched answers tailored to diverse user needs.
  • Prompt 4: This prompt was designed to position the model as an advisor trusted by farmers and policymakers alike, having scientific expertise with leadership, so that it can provide context-rich, evidence-based answers rooted in both practical farming knowledge and awareness of modern agricultural advancements.
  • Prompt 5: This prompt was kept minimal to evaluate the model’s inherent capabilities without heavy role conditioning, by providing only a brief instruction focused on expected output quality.
@@
Appendix C Human Evaluation Framework
This appendix details the evaluation framework used to assess the quality of LLM-generated answers in the agricultural domain. The evaluation is conducted by agriculture experts based on the following three criteria:
@@ 3. Factuality
Appendix D Contributors to Evaluation Dataset
The evaluation dataset used in this study for English, Hindi and Punjabi languages was curated by Kashish Mahendra, Vidhi and Sudhir Rajput. Our approach follows the principles of participatory research, and data creators have been offered coauthorship to compensate for their contributions.
Appendix E Hyperparameters
Parameter Value
Base Model LLaMA-3-8B-Instruct
LoRA Rank ($r$) 16
LoRA Alpha 16
Target Modules q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, down_proj
Batch Size 2
Gradient Accumulation Steps 4
Max Sequence Length 2048
Learning Rate 2e-4
Epochs 1
Optimizer AdamW (8-bit)
Quantization 4-bit (bnb)
Checkpointing Enabled (Unsloth)
Scheduler Linear with warm-up
Tracking Weights & Biases
Tokenizer LLaMA 3 Instruct tokenizer
Parameter Value
Base Model LLaMA-3.1-8B-Instruct
LoRA Rank ($r$) 16
LoRA Alpha 16
Target Modules q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, down_proj
Batch Size 2
Gradient Accumulation Steps 4
Max Sequence Length 2048
Learning Rate 2e-4
Epochs 1
Optimizer AdamW (8-bit)
Quantization Not Used
Checkpointing Enabled (Unsloth)
Scheduler Linear with warm-up
Tracking Weights & Biases
Tokenizer LLaMA 3 Instruct tokenizer
Appendix F Implementation Details
We employed Parameter-Efficient Fine-Tuning (PEFT) using the QLoRA method Dettmers et al. (2023) for adapting both the LLaMA-3-8B-Instruct and LLaMA-3.1-8B-Instruct models. PEFT was implemented via the Unsloth framework, which enables memory-efficient training through 4-bit quantization (used only for LLaMA-3) and Low-Rank Adaptation (LoRA). Gradient checkpointing was enabled to reduce memory usage during backpropagation.
Appendix G Hardware and Infrastructure
All experiments were conducted on high-performance GPU nodes equipped with NVIDIA Tesla V100 (32GB) and NVIDIA A100 (80GB) GPUs. The environment was configured to efficiently support large-scale language model inference and fine-tuning, leveraging mixed precision and distributed processing where applicable.
Appendix H Inter-annotator Agreement Calculation
@@
Appendix I Model Temperature Selection
Model T = 0.2 T = 0.5 T = 0.7
@@ Appendix I Model Temperature Selection
Table 13: English Model evaluation across 3 different temperature ($T$) settings using metrics: Agriculture Consensus (Con), Factuality (Fact), and Relevancy (Rel). The highest percentage is in bold.
Appendix J LLM-as-a-Judge : Pearson Correlation for Agriculture Consensus Metric
For the agriculture consensus metric, which involves binary classification, Pearson correlation Pearson et al. (1895); Fu et al. (2023) was considered. However, the correlation yielded a NaN value due to the absence of variance in the responses: both the annotators and the LLM provided identical labels for all instances, indicating complete agreement.
Appendix K LLM as Judge Prompts

### v2

@@ Appendix D Prompt Selection
Prompt Intuition
@@
  • Prompt 2: This prompt was designed to simulate a highly knowledgeable agricultural expert, embedding deep domain-specific authority into the LLM’s persona. The aim was to guide the LLM to generate precise, well-informed, and context-aware answers like domain experts.
  • Prompt 3: This prompt was designed to extract detailed, accurate, and context-sensitive responses by positioning the LLM as a multilingual agricultural expert. Emphasis was placed on reasoning, cultural relevance, and linguistic alignment to ensure comprehensive, well-researched answers tailored to diverse user needs.
  • Prompt 4: This prompt was designed to position the LLM as an advisor trusted by farmers and policymakers alike, having scientific expertise with leadership, so that it can provide context-rich, evidence-based answers rooted in both practical farming knowledge and awareness of modern agricultural advancements.
  • Prompt 5: This prompt was kept minimal to evaluate the LLM’s inherent capabilities without heavy role conditioning, by providing only a brief instruction focused on expected output quality.
@@
Appendix E Hyperparameters and Implementation Details
Parameter Value
Base LLM Llama-3-8B-Instruct
LoRA Rank ($r$) 16
LoRA Alpha 16
Target Modules
q_proj, k_proj, v_proj, o_proj,
gate_proj, up_proj, down_proj
Batch Size 2
Gradient Accumulation Steps 4
Max Sequence Length 2048
Learning Rate 2e-4
Epochs 1
Optimizer AdamW (8-bit)
Quantization 4-bit (bnb)
Checkpointing Enabled (Unsloth)
Scheduler Linear with warm-up
Tracking Weights & Biases
Tokenizer Llama 3 Instruct tokenizer
Parameter Value
Base LLM Llama-3.1-8B-Instruct
LoRA Rank ($r$) 16
LoRA Alpha 16
Target Modules
q_proj, k_proj, v_proj, o_proj,
gate_proj, up_proj, down_proj
Batch Size 2
Gradient Accumulation Steps 4
Max Sequence Length 2048
Learning Rate 2e-4
Epochs 1
Optimizer AdamW (8-bit)
Quantization Not Used
Checkpointing Enabled (Unsloth)
Scheduler Linear with warm-up
Tracking Weights & Biases
Tokenizer Llama 3 Instruct tokenizer
We employed Parameter-Efficient Fine-Tuning (PEFT) using the QLoRA method Dettmers et al. (2023) for adapting both the Llama-3-8B-Instruct and Llama-3.1-8B-Instruct LLMs. PEFT was implemented via the Unsloth framework, which enables memory-efficient training through 4-bit quantization (used only for Llama-3) and Low-Rank Adaptation (LoRA). Gradient checkpointing was enabled to reduce memory usage during backpropagation.
Appendix F Human Evaluation Framework
This appendix details the evaluation framework used to assess the quality of LLM-generated answers in the agricultural domain. Evaluators are native speakers of Hindi or Punjabi and are also fluent in English. The evaluation is conducted following three criteria:
@@ 3. Factuality
Appendix G Inter-annotator Agreement Calculation
@@
Appendix H LLM, Temperature and Prompt Selection
Prompt Number Consensus Relevancy Factuality
Prompt 1 100.0% 95.0% 55.0%
Prompt 2 100.0% 80.0% 50.0%
Prompt 3 100.0% 85.0% 45.0%
Prompt 4 100.0% 85.0% 50.0%
Prompt 5 100.0% 100.0% 40.0%
Prompt 6 100.0% 75.0% 40.0%
Table 15: Results for the prompts for English.
LLM T = 0.2 T = 0.5 T = 0.7
@@ Appendix H LLM, Temperature and Prompt Selection
Table 16: English LLM evaluation across 3 different temperature ($T$) settings using metrics: Agriculture Consensus (Con), Factuality (Fact), and Relevancy (Rel). The highest percentage is in bold.
Appendix I LLM as Judge Prompts
@@
Appendix J Hardware and Infrastructure
All experiments were conducted on high-performance GPU nodes equipped with NVIDIA Tesla V100 (32GB) and NVIDIA A100 (80GB) GPUs. The environment was configured to efficiently support large-scale LLM inference and fine-tuning, leveraging mixed precision and distributed processing where applicable. In total, approximately 180 GPU hours were utilized on the A100 (80GB) nodes for LLM fine-tuning and evaluation.
Experiment Description Hours Minutes
Fine-tuning
 Punjabi Llama3.1 Finetuning 3 12
 Hindi Llama3.1 Finetuning 2 17
 English Llama3 Finetuning 2 28
Model and Temperature Selection
 Airavata 12 03
 DeepSeek 29 06
 Llama3 17 36
 Mistral 12 09
Fine-tuning Inference
 English 1 46
 Hindi 2 36
 Punjabi 4 55
Individual Model Evaluations
 Llama3 (HI) 4 41
 Llama3.1 (HI) 7 54
 Gemma (HI) 1 59
 Mistral (HI) 6 11
 Navrasa (HI) 4 36
 Punjabi (HI) 4 55
 Llama3 (PA) 5 18
 Llama3.1 (PA) 6 21
 Navrasa (PA) 1 47
Prompt Selection Experiments
 Prompt 1 5 55
 Prompt 2 6 02
 Prompt 3 6 44
 Prompt 4 10 40
 Prompt 5 11 43
 Prompt 6 8 04
Indic Translation 5 48
Total GPU Time 178 46
Table 17: Detailed breakdown of GPU time consumed across all experiments including fine-tuning, inference, model selection, and prompt evaluation.
Appendix K LLM-as-a-Judge : Pearson Correlation for Agriculture Consensus Metric
For the agriculture consensus metric, which involves binary classification, Pearson correlation Pearson et al. (1895); Fu et al. (2023) was considered. However, the correlation yielded a NaN value due to the absence of variance in the evaluations: both the annotators and the LLM provided identical labels for all instances, indicating complete agreement.
Appendix L Misleading Score
Question Gold Answer LLM Answer Human LLM BERT ROUGE-L BLEU
Is pseudostem split symptom caused by wilt pathogens? Yes, the FOC can cause pseudostem split symptoms. Similar symptoms are also noticed with Erwinia rot disease. No, pseudostem split symptom is not caused by wilt pathogens. It is caused by the infestation of the stem borer, Chilo partellus. Incorrect Incorrect 0.650 0.153 0.030
Is pseudostem split symptom caused by wilt pathogens? Yes, the FOC can cause pseudostem split symptoms. Similar symptoms are also noticed with Erwinia rot disease. Yes, it is caused by wilt pathogens. Correct Correct 0.666 0.166 0.007
Can I apply the entire fertilizer recommended at one time for coconut plantation? Apply in split doses to ensure regular nutrient availability and avoid leaching. No, apply in controlled manner over time for optimal growth. Correct Partially Correct 0.621 0.170 0.019
Can I apply the entire fertilizer recommended at one time for coconut plantation? Apply in split doses to ensure regular nutrient availability and avoid leaching. Yes, you can apply the entire fertilizer recommended at one time. Incorrect Incorrect 0.567 0.091 0.000
Table 18: Evaluation of LLM-generated answers against gold references with human and LLM annotations along with automatic scores.